• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 19:38:28 on July30,2003

    Entry number 108221

    keyword=Magnets settled to new kinematics

    The HRS magnets are now settled to the next kinematic setting.

    Still waiting for the septum to cool down again.

    Doing Temperature tests (NMR) of the target in the meantime.

    FIGURE 1